If a 5kg body falls to the ground from a height of 10 meters and if all its mechanical energy is converted to heat the heat produced will be?

Grade:11

1 Answers

Kondala Goutham
6 years ago
Given mass of the body =5 kgIt falls to the ground from a height 10m,then the energy possessed by it =mgh=5×10×10=500 JThe mechanical energy possessed by the body =PE+KE=500 J+0 J=500 JThis energy is converted into heat energy=500 J